2/3 divided by u = 10/9 i can get the answer which is 20/27 explain how to get to that point

Answer by edjones(8007)

2/3/u = 10/9 I hope this is the equation.
2/3u=10/9 because (2/3)*(1/u)=2/3u
6=10u multiply 9u times both sides to get rid of the fractions.
u=6/10 divide each side by 10.
u=3/5 Ans.
Ed